Technical Staff Blog

Archives 2015

Last update on .

filemaker.jpeg We've been informed by CIS that

FileMaker A will be unavailable on 12/3/2015 at 6:00 AM to 8:00 AM. During this time, you will not be able to use access to databases hosted on Brown's FileMaker A server, either using FileMaker or via a web site.

If you experience ...

Last update on .

Eiki LC-WUL100A.jpg The projector in CIT 316 is experiencing some hardware issues and should be considered unreliable. We have ordered a replacement, but not yet have an ETA for arrival. If you have this room booked in the next week or so and are in need of projection, we encourage you to reach out to CS reception ...

Last update on .

ibm-gpfs.jpg One of the dept cluster nfs server, peeps, had a file locking problem around 13:45pm today causing a brief nfs blip for those linux machines connected to the 10.116.110.150 nfs ip address.  The locking issue on peeps has been resolved around 14:06pm.  Please email problem@cs.brown.edu if your ...

Last update on .

power-outage.jpg (09:29) Facilities has determined the root cause of yesterday's power outage is water seeping into electrical transfomers in the SciLi. The will be conducting emergency fixes starting @ 10am this morning that will necessitate taking down all power to the SciLi. This means our students currently assigned to space in the SciLi will need ...

Last update on .

power-outage.jpg No doubt you've heard by now, the CIT experiened a building wide power outage. This is the first time since we deployed our UPS that we've lost power longer than the batteries could hold us, so the entire machine room was powered down. As of 12:23pm, we believe most critical services are ...

Last update on .

network-2400px.png (10:35) Around 9:15am this morning, a network card in a CIS switch failed that connects our primary firewall's external network interface. This necessitated someone going onsite to manually force our firewalls to failover. Service was restored around 10:30am. We are in the process of checking various services to make sure that ...

Last update on .

logo_sympa_small.png At 07:38 this morning we discovered a dead process on our list server. Upon restarting the process a flurry of delayed emails were sent. We are investigating the root cause and, perhaps more importantly, why automated monitoring did not properly detect and correct the dead process.

Last update on .

xen_project_logo_dualcolor_165x69.png One of our Xen hosts seems to have had a disk failure over the past 12hrs, which resulted in nearly all of it's virtual machines going offline.  We have brought this machine back online temporarily while we work to temporarily move those virtual machines to a new host. 

The affected virtual machines are:

anvil ...

Last update on .

All shared network printers in the department have moved to the restricted service network in order protect them from printer attacks from inside the department.   What this means is direct connection to the printers are denied from inside the department or over the CS VPN.  All shared network printer connections must be mapped through the ...

Monthly archives

Next year

2016